OpenStudy (anonymous):

According to this property, 0n = 0. According to this property, 0n = 0. @Mathematics

OpenStudy (anonymous):

i am not sure what you want for this, but actually it is proved by the distributive property

OpenStudy (anonymous):

\[0n=(0+0)n=0n+0n\implies 0n = 0\]

OpenStudy (anonymous):

hello all!

myininaya (myininaya):

hey!

OpenStudy (anonymous):

@myininaya you missed some doozies last night. lots of wrong answers and not just typos from me

OpenStudy (turingtest):

I'm pretty sure you just want to know this is called the "zero product property". Will just the name of the rule suffice? Hello!

OpenStudy (anonymous):

what is "zero product property"? i am not saying you are wrong, i just have not heard of it

OpenStudy (anonymous):

i actually looked it up it says \[ab=0\implies a = 0 \text{ or } b=0\]

OpenStudy (turingtest):

I've heard it also called the identity property of zero. It's just the rule that states (0)x=0 no matter what x is. Her'es a lame link just to show I didn't make it up. http://www.sparknotes.com/math/prealgebra/operations/section3.rhtml

myininaya (myininaya):

0x=0 why does the name of the property matter?

OpenStudy (turingtest):

right? I agree, but they like to use it for formal proofs

OpenStudy (anonymous):

2myinninaya why is it true?

myininaya (myininaya):

i like to say this is true since 0x=0 in a proof

OpenStudy (anonymous):

0 is the identity of addition, so how do we know how it acts with multiplication

myininaya (myininaya):

i see nothing wrong with calling it the zero property

OpenStudy (anonymous):

and don't tell me "because anything times 0 is 0"!

OpenStudy (turingtest):

Is this part of Peano's axioms? I don't think so...

myininaya (myininaya):

a rectangle of height 0 with any base has area 0

OpenStudy (anonymous):

you can call it anything you like. i call myself satellite, and you call yourself "myininaya" or "turingtest" that is not my point

OpenStudy (turingtest):

Logically this is obvious, but if it's not an axiom it has to be proven by SOMEone

myininaya (myininaya):

\[0n=(k-k)n=nk-nk=0\]

OpenStudy (anonymous):

i proved it above assuming that multiplication distributes over addition. the distributive law is the link between times and plus, so that is what you have to use

OpenStudy (anonymous):

that looks like a proof as well. note the use of the distributive law!

OpenStudy (turingtest):

Never seen it proved before. Interesting. Later guys I have to go learn Spanish...

myininaya (myininaya):

yes i see the use of distributive law